# Graphical lasso for extremes

### v2

@@ 5.4.2 Diamond graph
We plot the values of $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ against the values of $M$ for parameters $x=0.8$ and $x=1.2$ on the lower two panels of Figure 2. Again we observe that while $\lambda_{2}$ is close the minimum of $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$, the interval $M\in[\lambda_{2},\lambda_{d}]$ in both examples satisfies the Mutual Incoherence condition.
Figure 2: The curves of $|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ versus $M$ for the star graph and the diamond graph, each with parameter values $x=0.8$ and $x=1.2$. The two dash vertical lines in each graph indicates the values of $\lambda_{2}$ and $\lambda_{d}$ for the corresponding $\Sigma$.
@@ 6 Simulations
We conduct a simulation study to show the performance of extreme graphical lasso algorithm. The simulation setup follows closely the one in Engelke et al. (2022), in order to maintain comparability.
@@ 6.1 Simulation setup
Throughout the simulation study, we sample random graph structures using a preferential attachment model with $d$ nodes and a degree parameter $q=1,2$, namely, the Barabasi–Albert model Albert and Barabási (2002), denoted as $BA(d,q)$. Note that $q=1$ corresponds to a tree model.
@@
For a given graph structure, we sample the precision matrix $\Theta$ as follows. For each edge in the graph, we simulate each non-zero upper triangular element in the matrix $\Theta$ from an independent uniform distribution $U[-5,-2]$. The elements in $\Theta$ corresponding to a non-edge is set to zero and the lower triangular elements are set to mirror their upper triangular counterparts. Finally, the diagonal elements in $\Theta$ are filled such that each row sum is zero.
We simulate $n$ observations from the multivariate Pareto distribution as defined in (2.2), where the spectral distribution function follows the HR model corresponding to the simulated precision matrix $\Theta$. Here we differ from Engelke et al. (2022) by using the multivariate Pareto distribution instead of the max-stable distribution. For each simulated sample, we follow the procedure in Section 5 to compute the sample covariance matrix $S$ using an intermediate sequence $k_{n}$ and then apply the extreme graphical lasso to estimate the precision matrix $\Theta$. The tuning parameter $M$ is chosen to be $M=\hat{\lambda}_{2}$, where $\hat{\lambda}_{2}$ is the estimated second largest eigenvalue of $S$, motivated by the theoretical guidance in Section 5.3. For the penalizing parameter $\gamma_{n}$, we tune it in a pre-specified range by minimizing a pseudo-(M)BIC metric defined in Section 6.2. For each setting, we repeat the simulations for $m=100$ samples. When evaluating the estimated graph, we use the F1-score defined as follows:
[…]
where $E$ and $\hat{E}$ are the edge sets for the true and estimated graphs, respectively.
6.2 Tuning the penalizing parameter $\gamma_{n}$
Recall from (5.1) that under the HR model,
$\left(\log\mathbf{Y}_{-k}-\log(Y_{k})\cdot\mathbf{1}\right)|_{Y_{k}>1}\sim N(\Gamma_{-k,k}/2,\Sigma^{(k)}).$
The negative log-likelihood based on a pseudo-likelihood can be written as
$-2\log L^{(k)}=k_{n}\left\{-\log|\Theta^{(k)}|+tr\left(\hat{\Sigma}^{(k)}\Theta^{(k)}\right)+(d-1)\log(2\pi)\right\},$
where $\Theta^{(k)}$ is the partial precision matrix to be estimated, $\hat{\Sigma}^{(k)}$ is calculated from (5.2) and $k_{n}$ is the total number of pseudo observations used in the estimation. We can then define a pseudo-MBIC metric as
$MBIC^{(k)}=k_{n}\left\{-\log|\Theta^{(k)}|+tr\left(\hat{\Sigma}^{(k)}\Theta^{(k)}\right)\right\}+p^{(k)}(\log\log(d-1))\log k_{n},$
where $p^{(k)}$ is the number of non-zero parameters in the non-diagonal part of $\Theta^{(k)}$ and we omit a constant term in the definition. While MBIC is also used in Engelke et al. (2022), in our case, we base our log-likelihood on the normal distribution instead and avoid evaluating the likelihood of an HR distribution. To maintain comparability, when evaluating a given estimated graphical structure, we re-estimate $\Theta$ using the original estimated $\hat{\Sigma}$ matrix and the matrix completion method as in Hentschel et al. (2022) based on the graph estimated by the extreme graphical lasso.
Given $MBIC^{(k)}$ for $1\leq k\leq d$, we can aggregate them into one metric by either taking the sum or the maximum. We remark that from Proposition 3.1, the sum approach is equivalent to a pseudo-MBIC metric defined for $\Theta$ directly. Empirically, both approaches choose the same optimal $\gamma_{n}$ in most cases. We opt to use the maximum approach
$maxMBIC=\max_{1\leq k\leq d}MBIC^{(k)}.$
and choose the penalizing parameter $\gamma_{n}$ that minimizes $maxMBIC$. In such a way, the corresponding estimated graphical structure ensures a goodness-of-fit for all subgraphs.
To demonstrate the tuning procedure, we conduct simulations for $d=20$. We simulate $m=100$ samples with $n=500$ from the $BA(d,q)$ model with $q=1,2$. In the estimation, we choose $k_{n}=75$ corresponding to top $15\%$ of observations. The pre-specified range of $\gamma_{n}$ is $\log_{10}(\gamma_{n})\in\{-2,-1.9,-1.8,\ldots,0\}$, containing 21 values in total. In Figure 3, we plot the average F1 score across these samples when using different values for the penalizing parameter. In addition, we plot the fraction of times when a specific value is chosen as the optimal penalizing parameter following the maxMBIC standard.
Figure 3: The average F1 score and the fraction of times being optimal versus the penalizing parameter $\gamma_{n}$ (in log scale) for the $BA(20,1)$ (upper panel) and $BA(20,2)$ (lower panel) graphs. The solid line provides the F1 score for the estimated graphs (left scale). The dash line provides the fraction of times (right scale) when each value is chosen as the optimal penalizing parameter. The plots are over 100 simulated samples.
For the $BA(20,2)$ model, we observe that the two lines peak at the same value for the penalizing parameter, demonstrating the usefulness of the tuning procedure. For the $BA(20,1)$ model, the most selected optimal values tend to be slightly lower than the penalizing parameter value leading to the highest F1 score. This could be due to the fact that the $BA(20,1)$ model corresponds to a tree structure. Given that the true graph is a tree, in estimation, a high penalizing parameter may results in an unconnected graph. Due to the evaluation of the log-likelihood, our $maxMBIC$ standard can only evaluate connected graphs. Therefore we stop the evaluation whenever a graph is unconnected. Consequently, for a more sparse graph such as a tree, the $maxMBIC$ tuning standard may result in a lower value for the chosen penalizing parameter for the sake of keeping the connectedness.
6.3 Graph discovery: large sample
To evaluate the graph recovery performance of the extreme graphical lasso, we perform a large sample simulation with $d=20$. More specifically, we simulate $m=100$ samples with $n=5000$ from the $BA(d,q)$ model with $q=1,2$. For each sample, we choose $k_{n}=500$ corresponding to top $10\%$ of observations to estimate the covariance matrix and then apply the extreme graphical lasso to estimate the graph structure. For tuning the penalizing parameter, we use the same range as in Section 6.2.
Figure 4 shows the aggregated graphs based on aggregating the 100 estimated graphs into one graph, where the thickness of each edge reflects the percentage among the 100 estimated graphs when this pair is identified as an edge. Comparing to the true graph, the extreme graphical lasso algorithm can often identify a true edge. Nevertheless, there are also “false positive” where an estimate edge is not an edge in the true graph. This is particularly the case for the $BA(d,1)$ model. Again, we attribute this to the fact that our algorithm aims at maintaining the connectedness of the graph, which often leads to under-penalization.
Figure 4: The aggregated graph structure based on 100 simulated samples. The true graph follows the $BA(20,1)$ (upper panel) and $BA(20,2)$ (lower panel) model.
6.4 Graph discovery: small sample
In this section, we evaluate the performance of the extreme graphical lasso when the sample size is relatively small relative to the dimension.
We simulate $m=100$ samples for both $d=20$ and $d=100$, from the $BA(d,q)$ models with $q=1,2$. For $q=1$, we consider three ratios $k_{n}/d\in\{1,2.5,5\}$. For $q=2$, we consider three ratios $k_{n}/d\in\{1,5,10\}$. Note that $k_{n}/d=1$ refers to a high dimensional scenario. We simulate $n=[k_{n}^{1/0.7}]$ observations for each given $k_{n}$, maintaining $k_{n}\approx n^{0.7}$. After estimating the covariance matrix $\Sigma$ , we apply the extreme graphical lasso to estimate the graph structure. For tuning the penalizing parameter, we use the same range as in Section 6.2 for $d=20$. To accelerate the calculation for $d=100$, we shrink the range of the tuning parameters to $\log_{10}(\gamma_{n})\in\{-2.1,-2,\cdots,-1.7\}$ containing five potential values.
Figure 5 shows the boxplots of the F1 score across each 100 sample for a given setup. For $d=20$, we obtain comparable F1 scores as that of the EGLearn algorithm with graphical lasso method, but worse than that of the EGLearn algorithm with neighborhood selection method, shown in Figure 5 in Engelke et al. (2022). In addition, the F1 scores are higher for higher $k_{n}/d$ ratio for the $BA(20,2)$ model. For the $BA(20,1)$ model, the impact of having a small sample is negligible. For $d=100$ and $q=2$, the F1 scores are comparable with that of the EGLearn algorithm with graphical lasso method as in Figure 6 in Engelke et al. (2022). The results for the $BA(100,1)$ model worsen. This could be again due to the fact that our tuning procedure is limited by maintaining connectedness. For both $q=1$ and $q=2$, we note the deterioration of performance when $k_{n}/d$ decreases.
Figure 5: The boxplots for the F1 scores versus different $k_{n}/d$ ratios, based on 100 simulated samples. The upper (lower) panel shows the result for $d=20$ ($d=100$). The left (right) graph is based on the $BA(20,1)$ ($BA(20,2)$) model.
7 Application
To illustrate the application of the extreme graphical lasso, we apply it to two real data examples.
The first dataset consists of spot foreign exchange rates denominated by the British Pound sterling between 2005 and 2020. The dataset consists of $n=3790$ daily observations with $d=26$ currencies, preprocessed by taking the absolute values of the de-garched log-returns, see Engelke and Volgushev (2022).
@@ 7 Application
Besides applying the extreme graphical lasso, we also apply the EGlearn with neighborhood selection algorithm as in Engelke et al. (2022). For both algorithms, we use $k_{n}=[n^{0.7}]$ for estimating the covariance matrix $\Sigma$ and the variogram $\Gamma$. For estimating the graphical structure, we use the same tuning procedure as in Section 6.1 and use the maxMBIC standard for the extreme graphical lasso. For the EGlearn algorithm, we follow the same tuning procedure as in Engelke et al. (2022).
Figure 6: Exchange rate data: identified graphical structures using the EGLearn and extreme graphical lasso (EGLasso) algorithms. The upper panels show the result based on the penalizing parameters turned by optimizing respective MBIC standards. The lower panels show sparser graphs using chosen penalizing parameters: $\rho=0.53$ for EGLearn, $\gamma_{n}=0.7$ for EGLasso.
The upper panels of Figure 6 show the estimated graphs from the two algorithms based on optimally tuned penalizing parameters. Both algorithms identify a large number of edges. In the lower panels of Figure 5, we present an estimated graph from each algorithm using a larger tuning parameter, resulting in sparser graphs that can be easier interpreted and compared. For the EGLearn algorithm, we follow Engelke et al. (2022) to choose $\rho=0.53$. For the extreme graphical lasso algorithm, we choose $\gamma_{n}=0.7$. Both algorithms reveal that there are two clusters among these currencies, one among Southern Asian countries, the other one among European countries. Observe from the EGLearn algorithm that the two clusters are connected via Japanese Yen, while the extreme graphical lasso results in an isolated node for Japanese Yen.
The second dataset concerns the river discharge at gauging stations from the upper Danube basin. This dataset, as preprocessed in Asadi et al. (2015), consists of $d=31$ stations and $n=428$ observations. We apply both the EGLearn with neighborhood selection algorithm and the extreme graphical lasso with $k_{n}=[n^{0.7}]$. For the EGLearn algorithm, we use $\rho=0.075$, $0.2$ and $0.405$, following the study of the same dataset in Engelke et al. (2022). For extreme graphical lasso, we use $\gamma_{n}=0.4$, $0.44$ and $0.445$.
Overall, as the penalizing parameters increase, the EGLearn algorithm inclines to produce more tree-like structure while the EGLasso algorithm produces more cluster structures with disconnected components. We remark that while the river network itself forms a tree network, the river discharge measurements can be influenced by factors other than the geographical locations of the gauging stations. For example, local rainfall or local snowmelt produces dependence in extreme river flow, as detected in the EGLasso algorithm.
Figure 7: Danube river flow data: identified graphical structures. Left column: the EGLearn algorithm with $\rho=0.075$, $0.2$ and $0.405$. Right column: extreme graphical lasso (EGLasso) algorithms with $\gamma_{n}=0.4$, $0.44$ and $0.445$.

### v3

@@ 5.4.2 Diamond graph
We plot the values of $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ against the values of $M$ for parameters $x=0.8$ and $x=1.2$ on the lower two panels of Figure 2. Again we observe that while $\lambda_{2}$ is close the minimum of $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$, the interval $M\in[\lambda_{2},\lambda_{d}]$ in both examples satisfies the Mutual Incoherence condition.
Figure 2: The curves of $|\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ versus $M$ for the star graph and the diamond graph, each with parameter values $x=0.8$ and $x=1.2$. The two dash vertical lines in each graph indicates the values of $\lambda_{2}$ and $\lambda_{d}$ for the corresponding $\Sigma$.
5.5 Correlation normalization
In practice, we normalize the estimated covariance matrix $S$ to a correlation matrix before applying the extreme graphical lasso. That is, we replace $S$ by $D^{-1/2}SD^{-1/2}$, where $D=\mathrm{diag}(S)$. This normalization ensures that the penalty parameter $\gamma_{n}$ has a consistent scale across different dimensions and different graph structures, simplifying the tuning procedure. After estimation, the precision matrix $\hat{\Theta}$ is transformed back to the original scale.
@@ 6 Simulations
We conduct a simulation study to evaluate the performance of the extreme graphical lasso (EGLasso) algorithm and compare it with the EGLearn algorithm with neighborhood selection (Engelke et al., 2022). Among the variants of EGLearn proposed in Engelke et al. (2022), the neighborhood selection variant outperforms graphical lasso as the base learner in every simulation scenario and has the strongest theoretical guarantees for consistent graph recovery.
@@ 6.1 Simulation setup
Throughout the simulation study, we sample random graph structures using the Barabási–Albert preferential attachment model (Albert and Barabási, 2002) with $d$ nodes and degree parameter $q=1,2$, denoted $BA(d,q)$. The case $q=1$ corresponds to a tree graph, while $q=2$ yields a denser graph.
@@
For a given graph structure, we sample the precision matrix $\Theta$ as follows. For each edge in the graph, we draw each non-zero upper triangular entry independently from a uniform distribution U[5,2]U[-5,-2]. The entries of $\Theta$ corresponding to non-edges are set to zero, the lower triangular entries mirror their upper triangular counterparts, and the diagonal entries are filled such that each row sum is zero. The resulting variogram matrix $\Gamma$ is then computed from $\Theta$.
We simulate $n$ observations from the max-stable Hüsler–Reiss distribution with variogram $\Gamma$. After simulating $n$ max-stable observations, we apply the peaks-over-threshold transformation: we retain the $k_{n}$ largest observations in terms of the radial component and transform them to approximate multivariate Pareto observations. Then, we compute the sample covariance matrix $S$ from the $k_{n}$ transformed exceedances and apply both the EGLasso and EGLearn algorithms. For the EGLasso, the tuning parameter $M$ is chosen to be $M=\hat{\lambda}_{2}$, where $\hat{\lambda}_{2}$ is the smallest positive eigenvalue of $S$, motivated by the theoretical guidance in Section 5.3.
To evaluate the estimated graph, we use the F1-score defined as follows:
$\text{F1-Score}=\frac{|E\bigcap\hat{E}|}{\frac{1}{2}\left(|E|+|\hat{E}|\right)}$,
where $E$ and $\hat{E}$ are the edge sets for the true and estimated graphs, respectively. For fair comparison, we report the oracle-tuned F1 score for both methods: for EGLasso, we select the penalty $\gamma_{n}$ that maximizes the F1 score over the grid $\log_{10}(\gamma_{n})\in\{-1.2,-1.1,\ldots,0\}$; for EGLearn, we analogously select the penalty $\rho$ that maximizes the F1 score over the grid $\rho\in\{0.175,0.2,0.225,\ldots,0.475\}$. Note for both algorithms, there are 13 values in the tuning grid. This isolates the effect of the estimation method from the tuning procedure. For each setting, we repeat the simulations for $m=100$ samples.
We maintain the sample size as $n=\lceil k_{n}^{1/0.7}\rceil$ so that $k_{n}\approx n^{0.7}$, consistent with the theoretical guidance for the intermediate sequence. We consider dimensions $d\in\{20,100\}$ with several $k_{n}/d$ ratios, resulting in a range of settings from the classical regime ($k_{n}/d=5$) to the high-dimensional regime ($k_{n}/d\leq 1$).
6.2 Graph recovery
Figure 3 displays boxplots of the F1 scores for both EGLasso and EGLearn across different settings. Both methods improve as the ratio $k_{n}/d$ increases, reflecting the benefit of having more exceedances relative to the dimension. For the denser $BA(d,2)$ graphs, EGLasso and EGLearn achieve comparable F1 scores: at $d=100$ and $k_{n}/d=5$, the median F1 scores are $0.87$ and $0.93$, respectively. For the tree graphs $BA(d,1)$, EGLearn with neighborhood selection has a clear advantage, particularly at $d=100$, where it achieves near-perfect recovery (median F1 $\approx 1.0$ at $k_{n}/d=2.5$) while EGLasso plateaus around $0.75$. This gap can be attributed to the fact that neighborhood selection exploits the conditional independence structure node-by-node, which is particularly effective for sparse tree graphs. In the high-dimensional regime $k_{n}/d\leq 1$, both methods struggle to achieve a high F1 score. Simulations based on multivariate Pareto samples (rather than max-stable samples) yield similar patterns. [^4] Results are available from the authors upon request.
Figure 3: Oracle-tuned F1 scores for EGLasso (blue) and EGLearn (green) under the max-stable data-generating process. Upper panel: $BA(d,1)$; lower panel: $BA(d,2)$. Left: $d=20$; right: $d=100$. Boxplots are over $100$ replications. The $x$-axis shows the ratio $k_{n}/d$.
6.3 Graph recovery: large sample illustration
To visualize the graph recovery of the EGLasso algorithm, we conduct a large-sample simulation with $d=20$ and $n=5000$, yielding $k_{n}=\lfloor n^{0.7}\rfloor=388$ exceedances ($k_{n}/d\approx 19$). For each of $m=100$ replications, we select the oracle penalty $\gamma_{n}$ that maximizes the F1 score.
Figure 4 displays the true graph alongside the aggregated estimated graph, where the thickness of each edge reflects the fraction of replications in which it is detected. True edges are shown in black and false positives in red. For the $BA(20,1)$ tree (upper panels), the EGLasso algorithm recovers the true edges consistently (median F1 =0.86=0.86 at median $\gamma_{n}=0.50$), with only a few false positive edges appearing at low frequency. For the denser $BA(20,2)$ graph (lower panels), the true edges are again detected with high frequency, though more false positives appear (median F1 =0.80=0.80 at median $\gamma_{n}=0.10$). The lower oracle $\gamma_{n}$ for $BA(20,2)$ reflects the denser true graph requiring less penalization to retain the larger number of true edges.
Figure 4: True graph (left) and aggregated estimated graph (right) based on 100 replications with $d=20$, $n=5000$. Upper panel: $BA(20,1)$ (tree); lower panel: $BA(20,2)$ (dense). Edge thickness in the aggregated graph reflects the detection frequency across replications. Black edges correspond to true edges; red edges are false positives. The oracle penalty $\gamma_{n}$ is selected per replication to maximize the F1 score.
6.4 Computational efficiency
Figure 5 displays boxplots of computation times (in seconds) for both methods. At $d=20$, both methods are fast (under one second per sample) and timing differences are negligible. At $d=100$, the computational advantage of EGLasso becomes apparent: for the $BA(100,2)$ graph, EGLasso is approximately 3355 times faster than EGLearn at $k_{n}/d\geq 1$, with median computation times of 5599 seconds compared to 23233131 seconds for EGLearn. For the $BA(100,1)$ tree, EGLasso achieves a 2244 times speedup at $k_{n}/d\geq 1$. At $k_{n}/d=0.5$, the rank deficiency of the sample covariance matrix $S$ leads to more iterations in the EGLasso algorithm, reducing the speed advantage.
The computational gain of EGLasso stems from its $O(d^{2})$ coordinate descent updates per iteration, compared to the $O(d^{3})$ matrix inversion steps required in each of the $d$ neighborhood regressions of EGLearn. Beyond computation time, EGLearn also has substantially higher memory requirements, as it stores $d$ separate regression problems simultaneously. To provide a fair comparison, all simulations were run on a computing cluster with 185GB RAM, ensuring that EGLearn was not limited by memory constraints. On a standard workstation with 64GB RAM, the speedup of EGLasso over EGLearn increases to approximately $17$ times for the $BA(100,2)$ setting at $k_{n}/d=1$. The gap in computation time will be further increased if multiple replications are run in parallel, because EGLearn processes will compete for limited memory.
Figure 5: Computation time (seconds) for EGLasso (blue) and EGLearn (green) under the max-stable data-generating process. Upper panel: $BA(d,1)$; lower panel: $BA(d,2)$. Left: $d=20$; right: $d=100$. Boxplots are over $100$ replications. The $x$-axis shows the ratio $k_{n}/d$.
7 Applications
@@
To illustrate the application of the extreme graphical lasso, we apply it to two real data examples. Besides applying the extreme graphical lasso, we also apply the EGLearn algorithm with neighborhood selection (Engelke et al., 2022). For each algorithm, we present estimated graphs at three penalty levels, with the penalty parameters chosen to yield a comparable number of edges. For both algorithms, we use $k_{n}=\lfloor n^{0.7}\rfloor$ for estimating the covariance matrix $\Sigma$ and the variogram $\Gamma$.
The first dataset consists of spot foreign exchange rates denominated by the British Pound sterling between 2005 and 2020. The dataset consists of $n=3790$ daily observations with $d=26$ currencies, preprocessed by taking the absolute values of the de-garched log-returns, see Engelke and Volgushev (2022). Figure 6 displays the estimated graphs for the exchange rate data. At $\gamma_{n}=0.35$ (62 edges), the EGLasso graph reveals a dense network with two main clusters: one among Southeast Asian currencies (HKG, TWN, MYS, THA, SGP, KOR) and another among European currencies (DNK, CZE, HUN, EUR, SWE, CHE). As the penalty increases to $\gamma_{n}=0.45$ (43 edges) and $\gamma_{n}=0.55$ (25 edges), these clusters become more pronounced and separated. The EGLearn algorithm at comparable edge counts ($\rho=0.25,0.35,0.55$) identifies a similar overall structure but tends to maintain connectivity as much as possible; for instance, at $\rho=0.55$ (27 edges), the EGLearn graph remains largely connected through bridge nodes, while EGLasso at $\gamma_{n}=0.55$ (25 edges) separates the Asian and European clusters more clearly.
Figure 6: Exchange rate data: estimated graphical structures using EGLearn (left) and EGLasso (right) at three penalty levels. The penalties are chosen to yield a comparable number of edges.
The second dataset concerns the river discharge at gauging stations from the upper Danube basin. This dataset, as preprocessed in Asadi et al. (2015), consists of $d=31$ stations and $n=428$ observations. Figure 7 shows the estimated graphs at three penalty levels. At $\gamma_{n}=0.45$ (88 edges) and $\gamma_{n}=0.55$ (68 edges), the EGLasso algorithm identifies clusters of stations that reflect local hydrological catchments, with several disconnected components. The EGLearn algorithm at comparable edge counts ($\rho=0.04$ and $0.07$) produces more connected, tree-like structures. At the sparsest level ($\gamma_{n}=0.70$, 53 edges vs $\rho=0.11$, 51 edges), both methods reveal a similar backbone structure following the river network.
Across both applications, a structural difference between the two methods is apparent: EGLearn with neighborhood selection tends to preserve global connectivity. By contrast, EGLasso penalizes the full precision matrix and can zero out all edges of a node simultaneously, producing disconnected components even at moderate sparsity levels. As a consequence, when the true underlying graph is expected to be connected (as in the river network), EGLearn captures the global topology more faithfully. On the other hand, when the goal is to discover local subgroups or clusters (as in the exchange rate data), EGLasso may yiled more informative subgroup structures.
Figure 7: Danube river flow data: estimated graphical structures using EGLearn (left) and EGLasso (right) at three penalty levels. The penalties are chosen to yield a comparable number of edges.